OpenStudy (anonymous):

PLEASE HELP!! Thank you!! :) There are 9 red gumballs, 5 green gumballs, 8 yellow gumballs, and 8 blue gumballs in a machine. Find P(green then yellow). A. 40/900 B. 5/30 C. 8/30 D. 40/870

OpenStudy (anonymous):

How do we find the probability of getting a green gum ball first?

OpenStudy (anonymous):

I think you find the sum of total gumballs in the machine, which is 30. Sorry I'm noot good at this :p

OpenStudy (anonymous):

most likely its B

OpenStudy (anonymous):

Right. Number of Green / Total Number = Probability of drawing a green first

OpenStudy (anonymous):

Now, after this, we've removed one green gumball. So now we only have 4 green in the machine, and 29 total in the machine. What's the probability of now drawing a yellow gum ball?

OpenStudy (anonymous):

8/29

OpenStudy (anonymous):

oh sorry I was wrong! lmao! please forgive me

OpenStudy (anonymous):

Good! Now, P(A then B) = P(A)*P(B) So we can calculate the probability of these two events happening one after the other.

OpenStudy (anonymous):

Lol it's fine @Xx_J.L.S_Xx Okay so P(5)*P(8)?

OpenStudy (anonymous):

Use the probabilities we found: P(Green first) = 5/30 P(Yellow second) = 8/29

OpenStudy (anonymous):

Oh, okay. Will brb gonna solve this. Lol

OpenStudy (anonymous):

Okay, I can't figure it out lol... Help?

OpenStudy (anonymous):

Watt nevermind! I got it it's 40/870 :)

OpenStudy (anonymous):

@Vandreigan he needs your help.

OpenStudy (anonymous):

oh ok :D good job!

OpenStudy (anonymous):

thanks for the medal!

OpenStudy (anonymous):

Thanks. Haha. Thanks for helping. :) thank you @Vandreigan :) you're welcome